OpenStudy (anonymous):

what is the substitution method for 9x + 5y = -44, -7x +y = 44

OpenStudy (anonymous):

replace y in the first equation by \[44+7x\]

OpenStudy (anonymous):

9x + 5y = -44 => x = (-44-5y)/9 substituting the x value in the either equation -7x +y = 44 =>-7[ (-44-5y)/9]+y=44 308+35y+9y=396 44y=88 y=2 =>x=(-44-10)/9=-6

OpenStudy (anonymous):

GOT IT!!!